OpenStudy (anonymous):

If f(θ) = cot θ, then f(π/2) is

OpenStudy (nincompoop):

\[\cot(\frac{ \pi }{ 2 })\]

OpenStudy (solomonzelman):

\[\large f(θ) = \cot θ\]\[\large f(\frac{π}{2}) = \cot (\frac{π}{2})\] \[\cot \frac{π}{2}=\cot{90}=0\]

OpenStudy (anonymous):

Thanks a lot

OpenStudy (solomonzelman):

Anytime !

OpenStudy (solomonzelman):

Also, for a future reference: you can find any section by: 1) clicking "more subjects" (the top left corner, towards the center) 2) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying). You can FAN someone by: 1) clicking on the user's name 2) clicking "fan" on the bottom right corner of user's box. ( you can also write a testimonial by clicking "add testimonial" and typing something in testimonial box) you can mention someone in your question (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @agnar2Lothbrok (they would get a notification that a user mentioned them in a reply) you can message someone by: 1) clicking on the someone's username 2) clicking "Send Message" 3) type your message press submit, and wait till the messages submits. If you have any questions you can message me or anyone else on this website. We will be happy to respond !

OpenStudy (anonymous):

thank you again

OpenStudy (solomonzelman):

My pleasure. This is a useful website for anyone who is on here... have a good luck !